OpenStudy (anonymous):

Average rate of change? f(x) =x^3 - 5x^2; x1 = 0, x2 = 2

OpenStudy (y2o2):

A(x) = ( f(2) - f(0) )/(2-0) = (2^3 -5(2²))/2 = -6